40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 23 00:05:17 2013, manasa, Update, SUS, Unstable Y arm ALS points to problems with ETMY suspension Y-SUS_All.png
    Reply  Wed Oct 23 09:17:33 2013, Steve, Update, SUS, ETMY sensors compared to ETMX ETMYglitching.pngOSEMsensorsETMY&X.pngETMYgettingKickedUp.pngdrive-damp.png
       Reply  Wed Oct 23 11:42:28 2013, rana, Update, SUS, ETMY sensors compared to ETMX 
       Reply  Wed Oct 23 15:48:06 2013, rana, Update, SUS, ETMY sensors compared to ETMX remote-control-cockroach.jpg
          Reply  Wed Oct 23 17:30:17 2013, jamie, Update, SUS, ETMY sensors compared to ETMX 
             Reply  Wed Oct 23 18:24:56 2013, rana, Update, SUS, ETMY sensors compared to ETMX 
                Reply  Wed Oct 23 18:28:01 2013, Jenne, Update, SUS, ETMY sensors compared to ETMX Screenshot-Untitled_Window.pngScreenshot-Untitled_Window-1.png
                   Reply  Thu Oct 24 10:31:25 2013, Steve, Update, SUS, ETMY sensors compared to ETMX 24h_glitching.pngdampingOffref.png
                   Reply  Tue Oct 29 10:51:16 2013, Masayuki, Update, SUS, ETMY sensors compared to ETMX 
                   Reply  Tue Oct 29 22:33:57 2013, rana, Update, SUS, ETMY sensors compared to ETMX Scratch-n-SniffStinkyDooBabyShowerGame1.jpg
                      Reply  Wed Oct 30 14:58:44 2013, Steve, HowTo, General, shutting down a computer 
                Reply  Thu Oct 24 12:00:11 2013, jamie, Update, CDS, fb acquisition of slow channels  
                   Reply  Fri Oct 25 01:46:33 2013, rana, Update, CDS, fb acquisition of slow channels  
Message ID: 9260     Entry time: Wed Oct 23 00:05:17 2013     Reply to this: 9262
Author: manasa 
Type: Update 
Category: SUS 
Subject: Unstable Y arm ALS points to problems with ETMY suspension 

[Masayuki, Manasa]

Looking into control signals and error signals of the Y arm green PDH servo,

1. The saturation of feedback signal (PZT_OUT) at +/-4000 counts (less than 5V) comes from only the readout saturating. The signals looked fine on the oscilloscope.
2. We did a sine sweep at the PZT_OUT and optimized the LO frequency. The LO frequency did not need any change.
3. The error signal has some offset to it. We are not sure where this comes from.

We have been seeing that whenever green loses lock, the spot position moves down in pitch on the ETMYF camera and the GTRY camera. This led us to think about if the lock loss originated from the PDH or from the cavity.

We looked into dataviewer channels of green, IR, oplev and suspension for the following cases:
1. Green and IR PDH locked
2. Green locked and arms flashing for IR
3. Green shutter closed and IR PDH locked
4. Green shutter closed and arms flashing for IR
5. Arms flashing for IR and ETMY oplev servo turned off.

Dataviewer snapshots of glitch in all the above cases are saved in masayuki's folder users/masayuki/ALS/kicked_mirror/

In all the above cases, we could still see the glitch. We could conclude that the problem lied with the ETMY SUS.
Shown below is the dataviewer snapshot of ETMY sus and shadow sensor channels. The glitch exists even when the oplev servo is turned off pointing to problems associated with the ETMY suspension.

Y-SUS_All.png

ELOG V3.1.3-